RC Update and Run Plan for March 22, 2020

From HallCWiki
Jump to: navigation, search

Current RC: Jay Benesch benesch@jlab.org

RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection

General updates:

Current Run Plan:

Punchlist for Sunday Day shift copied from URL below

https://hallcweb.jlab.org/wiki/index.php/PolHe3_Target_Coil_Rotation_Plan,_Mar_2020

Techs w/ RadCon [Todd (or designate) for cooling jets]

  • Turn on turbos for downstream beamline
  • Turn on turbos for SHMS and HMS spectrometer vacuums
  • Remove the guards on the SHMS and HMS entrance windows
  • Remove the guards on the both upstream and downstream beamline windows
    • RadCon will remove the bag on the upstream window
  • Ensure the beamline window cooling jets are in position and on
  • Ensure the target window cooling jets are in position and on
  • Install the target enclosure panels
  •  ??

Miscellaneous

  • [Gaskell] Beamline checklist
  • [OPS, Jay] Schedule hot checkout of Hall C arc magnets/power-supplies
  • [OPS, Jay] Verify all relevant FSDs that may have been masked for the down are known about and ready to unmasked when appropriate
  • [Lassiter] Enable all spectrometer magnet supplies (if not ready now?)
  • [Brad] Double check that rotation is enabled in the GUIs for both spectrometers (should be)
  • [Brad] Ramp all SHMS and HMS magnets to something modest (quick check), or to first kinematic. For example:
    • SHMS: 11.0deg, 7.5 GeV/c
    • HMS: 13.5deg, 4.2 GeV/c
    • NOTE: Coordinate with the target group on the appropriate HB setting so we don't interfere with any target pumping.
  • [Techs, Brad?] Do a final check to ensure that there are no potential issues with remote spectrometer rotation
  • Restart main alarm handler (resets masked channels)

- ??

Target Group

  • The camera in the laser box needs to be realigned
  • The rotating motor controller not fully working.
    • Needs Ethan’s help at some point to look into it
  • Transverse laser spectra-analyzer signal currently is not separate from longitudinal. We had it disconnected to avoid interference. May need to find a way to get separate signal.

Saturday-Sunday

"Shift Leader must request a beam that is shifted horizontally to the beam left side by 2 mm to prevent scraping."


Run Plan for Beam: Sunday SWING — Tuesday OWL

High level plan

  • Beam checkout (Call Brad when you think we are close)
  • Verify Cerenkov threshold in ELREAL trigger is set correctly
  • Take Optics data at our production kinematics (below)
    • Take data at the settings in the table below
      • NOTE: The times are Beam On Target . Do your best to estimate good quality 'beam on' time (ie. do not count trips or bad beam periods)
      • NOTE: The beam charge asymmetry does not matter for the optics runs!
    • Take and NMR every 2 hours to monitor spin up
  • Use maximum beam current that the Accelerator can provide and remain stable (up to a maximum of 45 uA on the Carbon target)
  • Set prescales to get close to 3.5--4 kHz accepted trigger rate in each arm
    • NOTE: If the un-prescaled trigger rates are low enough, then use 3/4 trigger instead of EL-REAL

Kinematic Settings

HMS SHMS TARGET POL I (uA) RASTER BEAM TIME Q (C) HMS TRIG SHMS TRIG STATUS COMMENT
P THETA P THETA (beam-on-target)
Optics: HMS KinA / SHMS KinX
Optics/No Sieve (Acceptance) 4.2 13.5 (-) 7.5 11 (-) C-Optics 40 4mm 0.5 0.05 3/4, ELREAL 3/4, ELREAL Done Verify ELREAL cer thr is correct
Verify ELCLEAN Cerenkov threshold is set correctly in both HMS and SHMS before proceeding.
Optics/No Sieve (Acceptance) 4.2 13.5 (-) 7.5 11 (-) C-Optics 40 4mm 0.5 0.05 ELREAL ELREAL Done Need >100k good electrons
No Target/No Sieve 4.2 13.5 (-) 7.5 11 (-) No Target 40 4mm 0.5 0.05 ELREAL ELREAL Need >100k good electrons
Optics/Sieve (angle & ytar) 4.2 13.5 (-) 7.5 11 (-) C-Optics 40 0.5mm 1 0.11 ELREAL ELREAL Done Need >50k good electrons
No Target/Sieve 4.2 13.5 (-) 7.5 11 (-) No Target 40 0.5mm 1 0.11 ELREAL ELREAL Need >50k good electrons
Optics: HMS KinC / SHMS KinZ
Optics/No Sieve (Acceptance) 4.2 16.4 (-) 6.4 14.5 (-) C-Optics 40 4mm 0.5 0.05 ELREAL ELREAL Need >100k good electrons
No Target/No Sieve 4.2 16.4 (-) 6.4 14.5 (-) No Target 40 4mm 0.5 0.05 ELREAL ELREAL Need >100k good electrons
Optics/Sieve (angle & ytar) 4.2 16.4 (-) 6.4 14.5 (-) C-Optics 40 0.5mm 1 0.11 ELREAL ELREAL Need >50k good electrons
No Target/Sieve 4.2 16.4 (-) 6.4 14.5 (-) No Target 40 0.5mm 1 0.11 ELREAL ELREAL Need >50k good electrons
Optics: HMS KinB / SHMS KinY
Optics/No Sieve (Acceptance) 4 20 (-) 5.6 18 (-) C-Optics 40 4mm 0.5 0.05 ELREAL ELREAL Need >100k good electrons
No Target/No Sieve 4 20 (-) 5.6 18 (-) No Target 40 4mm 0.5 0.05 ELREAL ELREAL Need >100k good electrons
Optics/Sieve (angle & ytar) 4 20 (-) 5.6 18 (-) C-Optics 40 0.5mm 1 0.11 ELREAL ELREAL Need >50k good electrons
No Target/Sieve 4 20 (-) 5.6 18 (-) No Target 40 0.5mm 1 0.11 ELREAL ELREAL Need >50k good electrons

Notes for the Future

  • At d2n startup, it would be a good idea to quantitatively check the dispersion at the target. I (D.Mack) don't think we need a tight spec, but if it's as sloppy as 1 cm/% (hence a decent spectrometer), then we're more at risk for things like how much the beam moves when the locks are off, or scraping from an off-energy tail on the beam. Jay:MQA3C13 can be adjusted to change horizontal dispersion.

Pending Issues

  • Be sure to restart main alarm handlers and turn up speaker volume before operating magnets! [1] -- Done
    Be sure to restart target alarm handler and turn up speaker volume before operating target! [2] -- Done
    Target Ladder software must throw an error on a 'voltage' readback mismatch! [3]
    Get 3H07B Harp repaired [4] [5]; and a puzzle [6]
    https://logbooks.jlab.org/entry/3807357 has Bill Henry's report on NMR and Kepco problems.
    https://logbooks.jlab.org/entry/3807176 has JP's report on laser polarization and alignment
    https://logbooks.jlab.org/entry/3807201 Has Arun's note about the unfortunate reference cell alignment. He implies
    NOTE: "Shift Leader must request a beam that is shifted horizontally to the beam left side by 2 mm to prevent scraping."
    pNMR issues
    Software 
    Software/driver issue needs to be resolved. Checkout during next non-production down time (Moller or SHMS fix)
    EPR problems 
    [7] saw some progress on Tuesday [8], but issues remain.
    Tentative plan now is to continue the d2n (and A1n) program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares).
    It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
    Kepco power supply 
    PS commissioning/installation (Todd, Bill) (Bill got a filter design from the vendor and has ordered parts.)
    New filter (for 180 Hz) to install for current mode operation fix. Can always go to voltage mode if it doesn't work.
    NMR Calibration Constants 
    Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted
    [9] [10] [11] [12] [13] (Bill, Junhao, Arun)
    Beam trips 
    Persistent BLM and ion chamber trips
    • General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
    • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
    • Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)
    USB bridge 
    Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
    • Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device
    • Ensure issues on Sat, March 14 have been corrected [14]
    iochc10 issues 
    Green now: Rebooted 3/2/20 with Dave Gaskell assistance. OPs wished to reboot hung iochc10 [15], noted in December as well [16]
    This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.
    Compton related items in Hall C beam steer-up procedure still cause confusion between MCC and Shift Crew

Opportunistic Access Jobs

None at the moment.



Bluejeans remote connection information

 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920


Useful Links